# Training LLMs over Neurally Compressed Text

### v2

@@ 6 Analysis
6.1 EqualInfoAC is less stable and less semantic than SentencePiece
@@
While the performance of our EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ model approaches that of our SentencePiece baseline, qualitative analysis shows that the two tokenization schemes differ in many regards. Table 7 illustrates some of these differences.
First, we observe that SentencePiece produces a stable text $\rightarrow$ token mapping. For example, “elephants” appears three times in the sentence, and maps stably to the same two-token sequence in all cases: [ elephant][s]. Similarly, both occurrences of “African” map to the same token: [ African]. By comparison, the EqualInfoAC tokenization is relatively unstable, with each occurrence of these words being segmented in a different way and yielding a different token sequence.
Second, we find that the SentencePiece tokenization is more “semantic”, by which we mean that the segmentation it induces aligns better with meaningful linguistic units—words and morphemes. While there are some exceptions, e.g. “savanna” being parsed as [s] [a] [v] [anna], the more common case is whole words being parsed as single tokens (e.g., currently), or into meaningful morphemes (e.g., elephant-s). By comparison, EqualInfoAC tokenization appears to almost entirely disregard word and morpheme boundaries. As one example, we see “Asian elephants.” parsed as [Asi] [an e] [lep] [hant] [s.].
@@ 6.2 AC decoding is learned step-by-step
To understand how a token’s position within a window affects learning, we track across training the average accuracy at each position within the $8$-token windows of a $403$m parameter EqualInfoAC$[b\mathord{=}64,\,v\mathord{=}256]$ model. [^21] The absolute accuracy of the EqualInfoAC$[b\mathord{=}64,\,v\mathord{=}256]$ model is relatively poor, but its relatively long window provides the clearest illustration of the positional trends. We observe similar trends for shorter windows where the model has stronger performance. Fig. 8 shows both raw accuracy (left) as well as the increase over “trivial” accuracy (right), which we define as the maximum accuracy achieved in the first $2{,}000$ steps of training. Looking at accuracy increase highlights the “sequential learning” trend by discounting any part of accuracy that is text independent. In particular, we note that window-final tokens have a non-uniform distribution due to the use of window-final padding bits (see our EqualInfoAC formulation in Section 3.3), which can be learned without any understanding of the text.
@@ Appendix C Instability in the Initial Token of Multi-Token Windows
Interestingly, (Külekci, 2011) point out this same issue, where a fixed size view of a variable length stream can cause false equivalencies when prefixes match. Similar to our findings, they find the models do have some limited ability to deal with these situations.
@@ D.3 Scaling Curves with Scaled Training Data
When scaling models, (Hoffmann et al., 2022) argue that training data should be scaled linearly with model size. As such, when comparing settings with constant training FLOPs, a large part of the FLOPs budget should be used by adding more training data. We apply this technique to compensate for our $2$b models being under-trained by plotting the scaling curves in Fig. 13, where the smaller models are trained with less data, proportional to their size. Models with $25$m parameters only train for $3$k steps, $113$m for $11$k, $403$m for $40$k, and $2$b for $200$k steps. Otherwise, the settings match those in Fig. 3. Numerical values used in the graph can be found in Table 13.
@@ D.6 Entropy Estimation
The maximum likelihood, or plug-in, estimator of entropy, $\hat{H}=-\sum_{x\in\mathcal{X}}\hat{p}(x)\log_{2}\hat{p}(x)$, is negatively biased—in fact, all entropy estimators are biased (Paninski, 2003). The Miller-Madow estimator attempts to correct for this bias by adding the approximate bias, caused by sampling, to the plug-in estimator. [^29] There are other methods for entropy bias correction such as (DeDeo et al., 2013) based on bootstrapping (Efron, 1979), however, with the size of the C4 training data, the required resampling was not possible. Thus, we use Miller-Madow in this work. The Miller-Madow estimator is given by $\hat{H}_{MM}=\hat{H}+\frac{\hat{|V|}-1}{2m}$. In this case, $m$ is the size of the sample used to estimate entropy and $\hat{|V|}$ is the estimated vocabulary size. In some applications, the vocabulary may often need to be estimated—for example new words may be added to languages—but in this case our vocabulary size is always $2^{n}$ where $n$ is the size of the current segmentation.

### v3

@@ 6 Analysis
6.1 AC-Based Tokenizations are Less Stable and Less Semantic than SentencePiece
Figure 8: Edit distance between tokenized sentence pairs that differ only by a single prefix word. With SentencePiece, changing the prefix has a minimal effect on tokenization, limited to the prefix itself. AC has the largest edit distances, as the prefix affects the probabilities M1 assigns to all subsequent text. EqualInfoAC falls in the middle, as many distinct prefixes lead to the same windowing of subsequent text, and each window is encoded in isolation.
We observe that text tokenized by our AC-based methods is not “stable”—that is, the tokenization of similar input data often leads to vastly different tokens sequences. To quantify this stability, we measure the effect of adding different prefixes to a fixed test sentence. We sample $500$ words from a list of $3{,}000$ common English words [^18] https://www.ef.com/ca/english-resources/english-vocabulary/top-3000-words/ and prepend each to the example sentence in Table 7. We then tokenize these new sentences using SentencePiece, AC$[v\mathord{=}256]$, and EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$. Finally, for each tokenizer, we calculate the Levenshtein edit distance (Levenshtein, 1965) between all pairs of tokenized sentences.
Fig. 8 shows a histogram of these distances. We observe that SentencePiece tokenized sentence pairs always have a low edit distance, as the effect of changing the prefix is limited to the prefix tokens. In contrast, the edit distance between AC-tokenized sentence pairs is consistently high. The presence of the prefix affects the probabilities M1 assigns throughout the entire sequence, resulting in widely varying token outputs. Edit distance with EqualInfoAC falls in between these two extremes. The degree to which the tokens after the prefix “match” depends on how the prefix aligns with EqualInfo window boundaries, which in turn affects the windowing of the rest of the sentence. Since there are a relatively small number of possible window alignments, many sentence pairs have highly overlapping token sequences. As AC tokenization is so much less stable than the other settings, we restrict further analysis to SentencePiece and EqualInfoAC.
@@ 6.1 AC-Based Tokenizations are Less Stable and Less Semantic than SentencePiece
While the performance of our EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ model approaches that of the SentencePiece baseline, our edit distances analysis suggests that the two tokenization schemes differ in many regards. To better understand these differences in qualitative terms, we compare the tokenizations of a single sentence in Table 7.
First, we observe that SentencePiece produces a stable text $\rightarrow$ token mapping. For example, “elephants” appears three times in the sentence, and stably maps to the same two-token sequence in all cases: [ elephant][s]. Similarly, both occurrences of “African” map to the same token: [ African]. In contrast, the EqualInfoAC tokenization is relatively unstable, with each occurrence of these words being segmented in a different way, and yielding different token sequences.
Second, we find that the SentencePiece tokenization is more “semantic”, by which we mean that the segmentation it induces aligns better with meaningful linguistic units—words and morphemes. While there are some exceptions, e.g. “savanna” being tokenized as [s] [a] [v] [anna], the more common case is that whole words are parsed as single tokens (e.g., currently), or into meaningful morphemes (e.g., elephant-s). By comparison, EqualInfoAC tokenization appears to almost entirely disregard word and morpheme boundaries. As one example, we see “Asian elephants.” tokenized as [Asi] [an e] [lep] [hant] [s.].
@@ 6.2 AC Decoding is Learned Step-by-Step
To understand how a token’s position within a window affects learning, we track the average accuracy at each position within the $8$-token windows of a $403$m parameter EqualInfoAC$[b\mathord{=}64,\,v\mathord{=}256]$ model [^22] The absolute accuracy of the EqualInfoAC$[b\mathord{=}64,\,v\mathord{=}256]$ model is relatively poor, but its relatively long window provides the clearest illustration of these positional trends. We observe similar trends for EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$ which has smaller windows but much stronger performance. during training. Fig. 10 shows both raw accuracy (left) as well as the increase over “trivial” accuracy (right), which we define as the maximum accuracy achieved in the first $2{,}000$ steps of training. By that point in training, the model predictions no longer change drastically between updates, and trivial patterns, such as how end-of-window padding generally reduces the number of possible tokens at the ends of windows, have already been learned. Looking at accuracy increase highlights the “sequential learning” trend by discounting any part of accuracy that is text independent. In particular, we note that window-final tokens have a non-uniform distribution due to the use of window-final padding bits (see our EqualInfoAC formulation in Section 3.3), which can be learned without any understanding of the text.
@@ Appendix C Instability in the Initial Token of Multi-Token Windows
Interestingly, Külekci (2011) point out this same issue, where a fixed size view of a variable length stream can cause false equivalencies when prefixes match. Similar to our findings, they find the models do have some limited ability to deal with these situations.
@@ D.3 Scaling Curves with Scaled Training Data
When scaling models, Hoffmann et al. (2022) argue that training data should be scaled linearly with model size. As such, when comparing settings with constant training FLOPs, a large part of the FLOPs budget should be used by adding more training data. We apply this technique to compensate for our $2$b models being under-trained by plotting the scaling curves in Fig. 15, where the smaller models are trained with less data, proportional to their size. Models with $25$m parameters only train for $3$k steps, $113$m for $11$k, $403$m for $40$k, and $2$b for $200$k steps. Otherwise, the settings match those in Fig. 3. Numerical values used in the graph can be found in Table 13.
@@ D.6 Entropy Estimation
The maximum likelihood, or plug-in, estimator of entropy, $\hat{H}=-\sum_{x\in\mathcal{X}}\hat{p}(x)\log_{2}\hat{p}(x)$, is negatively biased—in fact, all entropy estimators are biased (Paninski, 2003). The Miller-Madow estimator attempts to correct for this bias by adding the approximate bias, caused by sampling, to the plug-in estimator. [^30] There are other methods for entropy bias correction such as those used in DeDeo et al. (2013) based on bootstrapping (Efron, 1979), however, with the size of the C4 training data, the required resampling was not possible. Thus, we use Miller-Madow in this work. The Miller-Madow estimator is given by $\hat{H}_{MM}=\hat{H}+\frac{\hat{|V|}-1}{2m}$. In this case, $m$ is the size of the sample used to estimate entropy and $\hat{|V|}$ is the estimated vocabulary size. In some applications, the vocabulary may need to be estimated—for example, to account for new words that are added to languages—but in this case our vocabulary size is always $2^{n}$ where $n$ is the size of the current segmentation.